Find ∫ π‘₯π‘π‘œπ‘  π‘₯𝑑x


1
Expert's answer
2021-06-15T05:01:00-0400

Solution.


"\int x \cos xdx"


To calculate this integral we use the method of integration by parts.

Let be "u=x, dv=\cos {x} dx."

From here "du=1, v=\sin x."

Therefore, "\int x\cos xdx=x\sin x-\int 1\cdot\sin xdx=x\sin x+\cos x+C,"

where "C" is some constant.

Answer.


"\int x\cos xdx=x\sin x+\cos x+C."